OpenStudy (anonymous):

In how many different ways can the top nine new indie bands be ranked on a top nine list? The top hit song for each of the nine bands will compete to receive monetary awards of $1000, $500, $250 and $100, respectively. In how many ways can the awards be given out?

OpenStudy (anonymous):

\[9\times 8\times 7\times 6\]

OpenStudy (anonymous):

that was part 2. part 1 is 9!

OpenStudy (siddharth):

There are 9 possible choices for rank 1, There are 8 remaining choices for rank 2, since one band has been chosen for rank 1 already. There are 7 remaining choices for rank 3, since 2 bands are chosen already.. ..... So, the total possible ways are 9x8x7x6.... x2x1 which is 9!